OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal... Please help me to simplify this Rational Expression - 5x x 7x --- + --- + --- 4 6 12

OpenStudy (michele_laino):

hint: what is the minimum common factor?

OpenStudy (abbycross167):

@karim728 could you help me? I know the LCM is 12 and I need to multiply the top by 2 right?

OpenStudy (michele_laino):

next you have to compute these quantities: \(12/4=...\) \(12/6=...\) \(12/12=...\)

OpenStudy (abbycross167):

so 12/4=3 12/6=2 12/12=1?

OpenStudy (michele_laino):

correct! so, you have to multiply the first numerator by \(3\), the second numerator by \(2\), and the third numerator by \(1\), what we get, is then: \[\frac{{5x}}{4} + \frac{x}{6} + \frac{{7x}}{{12}} = \frac{{\left\{ {\left( {5x} \right) \cdot 3} \right\} + \left( {x \cdot 2} \right) + 7x}}{{12}} = ...?\] please simplify

OpenStudy (abbycross167):

wait so turn this {(5x)*3} + (x*2)+7x ----------------- 12 into this (15x)+(x^2)+7x -------------- 12

OpenStudy (michele_laino):

we have \(2 \cdot x= 2x\)

OpenStudy (michele_laino):

namely: \[\begin{gathered} \frac{{5x}}{4} + \frac{x}{6} + \frac{{7x}}{{12}} = \frac{{\left\{ {\left( {5x} \right) \cdot 3} \right\} + \left( {x \cdot 2} \right) + 7x}}{{12}} = \hfill \\ \hfill \\ = \frac{{15x + 2x + 7x}}{{12}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so, what is: \[{15x + 2x + 7x}=... \]

OpenStudy (abbycross167):

24x?

OpenStudy (michele_laino):

correct! and what is: \[\frac{{24}}{{12}} = ...?\]

OpenStudy (abbycross167):

1/2?

OpenStudy (michele_laino):

are you sure? \(24/12=...?\)

OpenStudy (michele_laino):

hint: I think that \(24/12=2\), am I right?

OpenStudy (abbycross167):

oh sorry, yes its 2

OpenStudy (michele_laino):

correct! So the final result is \(2x\): \[\begin{gathered} \frac{{5x}}{4} + \frac{x}{6} + \frac{{7x}}{{12}} = \frac{{\left\{ {\left( {5x} \right) \cdot 3} \right\} + \left( {x \cdot 2} \right) + 7x}}{{12}} = \hfill \\ \hfill \\ = \frac{{15x + 2x + 7x}}{{12}} = 2x \hfill \\ \end{gathered} \]

OpenStudy (abbycross167):

Awesome!!! thank you so much!!!

OpenStudy (michele_laino):

:)